OpenStudy (anonymous):

divide and simplify 7/25 (divided by) 4/5

OpenStudy (anonymous):

\[\frac{\frac{7}{25}}{\frac{4}{5}}\] \[=\frac{7}{25}*\frac{5}{4}\] \[\frac{7}{5}*\frac{1}{4}\] \[\frac{7}{20}\]